# A reanalysis of the FDA’s benefit-risk assessment of Moderna’s mRNA-1273 COVID vaccine stratified not only based on age and sex but also on prior-infection and comorbidity status

### v3

@@ 4.1. Subpopulation Analyses Based on Prior-Infection Status
There is a compelling case for conducting a benefit-risk assessment for the subpopulation of unvaccinated 18–25-year-old males having prior-infection protection because prior infection not only decreases the hospitalizations-prevented benefit of vaccination but also increases the risk of severe adverse events101010e.g., those described in [57, 58, 59, 60, 61, 62, 63] after vaccination [64, 65], with [64] reporting prior infection is “associated with an increased risk of severe side effects leading to hospital care (1.56 [risk ratio] (95%CI: 1.14–2.12)).” Moreover, prior infection may increase the risk of VAM/P [42]—see Appendix 4 of Supplement 2 for details. [^11] Survey results reported in [66] also suggest that prior infection may increase the risk of VAM/P; e.g., in Table 4, we see that 2.53% of those having a confirmed SARS-CoV-2 infection before dose 1 of mRNA-1273 reported “chest tightness/pain” during the 7-day period after the dose while only 0.54% of those not having a confirmed infection before dose 1 reported these symptoms; corresponding percentages for dose 2 (Table 5) are 1.78% and 0.98%.
For the subpopulation of 18–25-year-old unvaccinated males having a COVID-infection prior to the evaluation period 1/1/22 – 5/31/22, we can use our model to compute the expected number of COVID hospitalizations in various scenarios simply by setting $F_{pi}=1$ while leaving all other inputs unchanged. For example, with $F_{pi}=1$ in Scenario A, our model predicts $U\approx 155.82$ hospitalizations per million and $V\approx 72.24$, yielding a hospitalizations-prevented benefit of 83.58 per million vaccinated, far below VAM/P hospitalization risk ($\approx$ 268 per million vaccinated).
As we discuss in Appendix 3 of Supplement 1, the preceding paragraph’s analysis may be reinterpreted in terms of evaluation-period hospitalization risk for a typical unvaccinated 18–25-year-old male, COVID-infected prior to the evaluation period. For such a male, the total hospitalization risk of the choice to vaccinate is essentially the sum of his probability of VAM/P hospitalization and his probability of having a breakthrough-case COVID hospitalization during the evaluation period. Our analysis is summarized in Tables 7 and 8 below in terms of hospitalization-risk ratios, with ratios <1<1 indicating the choice to vaccinate decreases risk and >1>1 indicating it increases risk.
@@
11 155.82 in 1 million 72.24 in 1 million $2.18\approx\frac{268+72.24}{155.82}$
$0.5$ 77.91 in 1 million 36.12 in 1 million $3.90\approx\frac{268+36.12}{77.91}$
$0.25$ 38.96 in 1 million 18.06 in 1 million $7.34\approx\frac{268+18.06}{38.955}$
@@
$I_{r}(1-E_{pi})$ is the probability the “typical male” (18–25 with prior-infection protection) will become COVID infected over the evaluation period if unvaccinated; e.g., 0.53 if $I_{r}=1$
According to Table 7, our modeling suggests that for a typical unvaccinated 18–25-year-old male with prior-infection, the choice to vaccinate created a hospitalization risk over twice that associated with the choice not to vaccinate even if $I_{r}=1$ (evaluation-period infection is certain for an infection-naïve male 18–25). Because a “typical male” has an “average comorbidities level,” the information in Table 7 is difficult to apply. However, if our typical unvaccinated male of Table 7 also has no comorbidities, then his hospitalization-risk ratios should exceed those of Table 7, [^12] See, e.g., the BMI 23–24 column of Table 9 because risk-ratio denominators should decrease given no comorbidities and there is no evidence that being free of comorbidities reduces VAM/P risk [67]. [^13] See also [68, p. 383].
@@
Table 8. Hospitalization-risk ratios for a typical male 18—25 without prior-infection protection: risk of hospitalization after choosing mRNA-1273 vaccination (VAM/P or severe COVID-breakthrough infection) divided by risk of COVID hospitalization after choosing not to vaccinate.
@@
4.2. Subpopulation Analyses Based on Prior-Infection Status and BMI
@@
The FDA acknowledges, “The health condition of individuals may remarkably impact the B-R profile and its evaluation” [11, Section 4]. Comorbidities such as obesity may triple COVID-hospitalization risk [69]. Our model can produce a benefit-risk analysis based on comorbidity status by appropriately adjusting its infection-hospitalization-rate input $H_{r}$. We illustrate the process in Appendix 3 of Supplement 1, focusing on obesity because evidence suggests that Body Mass Index (BMI) is the single most potent modifier of COVID-19 risk for those under 60 years old [70, 71]. We use an assessment of COVID-hospitalization risk based on BMI that finds risk is minimized when BMI $\approx 23$ and that for $23\leq\text{BMI}\leq 44$ risk is multiplied by a factor of approximately 1.09 (1.08, 1.10) for each unit increase in BMI [72]. The results of our modeling appear in Table 9 below.
@@
Consistent with Tables 7 and 8, Table 9 indicates that for those having BMI in the intermediate range 25–34, prior-infection status determines whether vaccination is beneficial. However, for the those with high BMI (35–44) in a high-infection-risk scenario Table 9 indicates that vaccination becomes beneficial even for those with prior-infection protection.
4.3. Validating the Model
@@
In Supplement 2, we use Ontario COVID-hospitalizations data from the evaluation period 1/1/22– 5/31/22 to show that our model’s projections of Table 6 for the hospitalizations-prevented benefit of vaccination for 18–25-year-old males are much more accurate than those of the FDA [11]. [^14] We use Ontario data for many reasons; e.g., unlike the CDC, Public Health Ontario chose to publish incidental COVID-hospitalization rates (starting 1/8/2022) [30]; see [Supplement 2, Section 3] for other reasons. We test our model in two ways: (i) showing that it fits Ontario COVID-hospitalization data with inputs at plausible levels for Ontario over the evaluation period [Supplement 2 Section 8], and (ii) using Ontario COVID-hospitalization data to predict corresponding data for the U.S. [Supplement 2, Section 7]. For (i), we fit our model to Ontario data for the general-population mix of 18–25-year-old male-and-female Ontarians by using Mathematica to solve a sequence of least-squares problems depending on the value of $H_{r}$. For $H_{r}=0.0011$ and $0.0010$, we obtain model inputs in plausible ranges for Ontario that yield a perfect modeling fit and note that these values of $H_{r}$ are consistent with COVID hospitalization data from the state of Connecticut. A brief discussion of (ii) is provided in the next paragraph.
@@
Based on COVID hospitalization counts from Public Health Ontario as well as corresponding incidental hospitalization rates, we estimate that vaccinating one million 18–25-year-old Ontarians prevented 152 COVID hospitalizations over the evaluation period—see Table S2.5. Making various adjustments e.g., accounting for the better health of Ontarians vs. Americans, we obtain an upper-bound estimate that 214 evaluation-period hospitalizations were prevented by one million full COVID vaccinations of 18–25–year-old males in the U.S.—see [Supplement 2, Table S2.7] and the discussion that follows it. Observe that 214 is an upper bound for our projections of Scenarios B–E (see Table 6) but not an upper bound for any of the FDA’s projections.
@@
Ontario data also supports the definitiveness of the results of Table 7 above given the male has no comorbidities, suggesting that 151 in 1 million is an overestimate of his evaluation-period COVID hospitalization risk [Supplement 2, Section 2.3].
@@ 5. Discussion
By accounting for prior-infection protection and using evidence-based model inputs, we find in our Scenarios B–E of Table 6 that risks of mRNA-1273 vaccination outweighed benefits for the general population of 18–25-year-old males, assuming hospitalizations are the principal endpoint of analysis. In particular, for our scenarios predicting plausible infection-level increases over the evaluation period (Scenarios B, D, and E), our model suggests that vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myo/pericarditis (VAM/P) alone compared to COVID hospitalizations prevented. [^15] The range 16%–63% is consistent with our model-validation work discussed in Section 4.3: recall we estimate that at most 214 evaluation-period hospitalizations were prevented by one million mRNA-1273 vaccinations of 18–25-year-old males in the U.S. [Supplement 2, Table S2.7], and note that our corresponding VAM/P hospitalization projection of 268 is a little over 25% higher than 214. Recalling that the FDA found, “[T]he benefits of vaccination with the Moderna COVID-19 vaccine clearly outweigh the risks, even among the male adolescent population” [11, Section 4], we have demonstrated that the conclusions of a benefit-risk assessment for vaccination may be dramatically impacted by accounting for the benefits conferred by prior infection. We have also demonstrated in Tables 79 how the benefit-risk picture for vaccination may vary based on an individual’s prior-infection and comorbidity status.
@@
Consistent with our model’s finding that for 18–25-year-old males in the U.S. the risks of mRNA-1273 vaccination outweighed the benefits relative to hospitalizations, we note that France suspended mRNA-1273 use on 10/15/2021, making it available again on 11/8/21 but at half-dose and only for those over 30 years [73, p. 4].
@@
We emphasize that the cardiac risk of COVID infection is incorporated into our modeling because our estimates of COVID hospitalizations include those for COVID-related cardiac issues. We address here assertions such as “Importantly, the risk of myocarditis after any dose of COVID-19 vaccine, in any population, is substantially lower than the risk of myocarditis after SARS-CoV-2 infection,” which appeared, without supporting references, in an April 2024 editorial of the European Heart Journal [74]. Contrary to the editorial’s assertion, Patone et al. find that “In men younger than 40 years old, the number of excess myocarditis events [leading to hospitalization] per million people was higher after a second dose of mRNA-1273 than after a positive SARS-CoV-2 test (97 [95% CI, 91–99] versus 16 [95% CI, 12–18])” [6]. Karlstad et al. also find significantly higher risk of hospitalized myocarditis after dose 2: 183.9 per million 2nd mRNA-1273 doses [5, Table 2] vs. 13.7 per million positive tests [5, eTable 7] among males 16–24. Moreover, these myocarditis incidence rates expressed in cases per million positive tests overstate the risk of myocarditis after COVID infection expressed in cases per million COVID infections (because many infections are not confirmed by positive tests, but essentially all infections among  those seeking care for myocarditis are confirmed by positive tests) [75].  Consistent with [75], the authors of [76] note, “[T]he attributable risk estimates for COVID-19 [associated myocarditis] used laboratory confirmed cases as the denominator and will be affected by the proportion of all SARS-CoV-2 infections captured by testing, precluding a direct comparison with vaccine-associated attributable risks.”
The strength of prior-infection protection
@@
As noted in Section 2.2, the FDA’s assumption that prior infection would provide no protection against Omicron is contrary to immunological expectations and studies (e.g., [15, 16, 17, 19, 26, 27]) available at the time of the FDA’s mRNA-1273 analysis suggesting that prior-infection protection would likely equal or exceed vaccine protection among the COVID-naïve. As noted in Table 5 footnotes, we have assumed that vaccination’s effectiveness against hospitalization is 77% while that provided by prior infection is 88.9%. A systematic review [24] involving 21 vaccine-effectiveness studies from the period 12/3/2021 – 4/7/2022 reports,
@@
In contrast, the meta-analysis [21] estimates prior-infection’s effectiveness against severe Omicron to be over 80% through 6 months after infection, while the meta-analysis [22, Table S2] estimates its effectiveness to be over 88% through 45 weeks after infection, with the former estimate based on studies from 1/1/22–6/1/22 and the latter, on studies published by 9/31/22. In mid-January of 2022, the minimum reasonable assumption the FDA might have made is that prior-infection protection would equal vaccine protection (as in our Scenario E); instead, in all Omicron-based scenarios of [11], the agency ascribed to vaccination 72% effective protection against hospitalization while assuming prior infection offers no protection.
5.1. Recommendations
Our modeling provides a call to action to policy-forming health agencies to employ multi-variable models like ours that can better stratify vaccination recommendations. To protect vaccine acceptance and promote public trust, the FDA must adopt a rigorous framework for benefit-risk assessment of new vaccines and revisit recommendations as new data becomes available. In particular, the framework must incorporate consideration of the protection that prior infection affords against development of severe disease.
We also recommend United States public-health agencies reassess their priorities. Two years after the first cases of COVID were identified in the U.S., the FDA decided to “not investigate the benefits and risks [of mRNA-1273 vaccination] for specific subpopulations, such as those with comorbidities, due to limited information for these populations” [11, Section 4]. Also, public-health agencies had reason to prioritize understanding the effectiveness of prior-infection protection and had evidence that it was likely to be high. For example, in Pfizer’s phase-3 trial for BNT162b2, 567 participants in the placebo group and 526 in the treatment arm had evidence of prior COVID-19 infection. In each of these cohorts, there was only 1 reinfection (or <0.2<0.2% for both) [77, p. 27].

### v4

@@ 3.1. Subpopulation Analyses Based on Prior-Infection Status
There is a compelling case for conducting a benefit–risk assessment for the subpopulation of unvaccinated 18–25-year-old males having prior-infection protection because prior infection not only decreases the hospitalizations-prevented benefit of vaccination but also increases the risk of severe adverse events (e.g., those described in [60, 61, 62, 63, 64, 65, 66]) after vaccination [67, 68], with [67] reporting prior infection is “associated with an increased risk of severe side effects leading to hospital care (1.56 [risk ratio] (95%CI: 1.14–2.12)).” Moreover, prior infection may increase the risk of VAM/P [15]—see Appendix S4 of Supplement S2 for details. [^13] Survey results in [69] also suggest that prior infection may increase VAM/P risk; e.g., 2.53% of those having a confirmed SARS-CoV-2 infection before dose 1 of mRNA-1273 reported “chest tightness/pain” during the 7-day postdose period [69, Table 4] while only 0.54% of those not having a confirmed infection before dose 1 reported these symptoms; corresponding percentages for dose 2 are 1.78% and 0.98% [69, Table 5].
For the subpopulation of 18–25-year-old unvaccinated males having a COVID infection prior to the evaluation period 1 January 2022–31 May 2022, we can use our model to compute the expected number of COVID hospitalizations in various scenarios simply by setting $F_{pi}=1$ while leaving all other inputs unchanged. For example, with $F_{pi}=1$ in Scenario A, our model predicts $U\approx 161.7$ hospitalizations per million and $V\approx 72.24$, yielding a hospitalizations-prevented benefit of 89.46 per million vaccinated, far below VAM/P-hospitalization risk ($\approx$268 per million vaccinated).
As we discuss in Appendix S3 of Supplement S1, the preceding paragraph’s analysis may be reinterpreted in terms of evaluation-period hospitalization risk for a typical unvaccinated 18–25-year-old male COVID-infected prior to the evaluation period. For such a male, the total hospitalization risk of the choice to vaccinate is essentially the sum of his probability of VAM/P hospitalization and his probability of having a breakthrough-case COVID hospitalization during the evaluation period. Our analysis is summarized in Tables 6 and 7 below in terms of hospitalization-risk ratios, with ratios <<1 indicating the choice to vaccinate decreases risk and >>1 indicating it increases risk.
@@
11 161.70 in 1 million 72.24 in 1 million $2.10\approx\frac{268+72.24}{161.70}$
$0.5$ 80.85 in 1 million 36.12 in 1 million $3.76\approx\frac{268+36.12}{80.85}$
$0.25$ 40.43 in 1 million 18.06 in 1 million $7.08\approx\frac{268+18.06}{40.43}$
@@
$I_{r}(1-E_{pi})$ is the probability the “typical male” (18–25 with prior-infection protection) will become COVID infected over the evaluation period if unvaccinated; e.g., 0.55 if $I_{r}=1$
According to Table 6, our modeling suggests that for a typical unvaccinated 18–25-year-old male with prior infection, the choice to vaccinate created a hospitalization risk over twice that associated with the choice not to vaccinate even if $I_{r}=1$ (evaluation-period infection is certain for an infection-naïve male 18–25). Because a “typical male” has an “average comorbidities level,” the information in Table 6 is difficult to apply. However, if our typical unvaccinated male of Table 6 also has no comorbidities, then his hospitalization risk ratios should exceed those of Table 6, [^14] Compare, e.g., the BMI 23–24 column of Table 8 to the final column in Table 6. because risk-ratio denominators should decrease given no comorbidities and there is no evidence that being free of comorbidities reduces VAM/P risk [70]. [^15] See also [71, p. 383].
@@
Table 7. Hospitalization-risk ratios for a typical male 18—25 without prior-infection protection: risk of hospitalization after choosing mRNA-1273 vaccination (VAM/P or severe COVID-breakthrough infection*) divided by risk of COVID hospitalization after choosing not to vaccinate.
@@
*COVID hospitalization risk is considered over the 5-month evaluation period.
3.2. Subpopulation Analyses Based on Prior Infection Status and BMI
@@
The FDA acknowledges, “The health condition of individuals may remarkably impact the B-R profile and its evaluation” [13, Section 4]. Comorbidities such as obesity may triple COVID-hospitalization risk [72]. Our model can produce a benefit–risk analysis based on comorbidity status by appropriately adjusting its infection-hospitalization-rate input $H_{r}$. We illustrate the process in Appendix S3 of Supplement S1, focusing on obesity because evidence suggests that Body Mass Index (BMI) is the single most potent modifier of COVID-19 risk for those under 60 years old [73, 74]. We use an assessment of COVID-hospitalization risk based on BMI that finds risk is minimized when BMI $\approx 23$ and that for $23\leq\text{BMI}\leq 44$ risk is multiplied by a factor of approximately 1.09 (1.08, 1.10) for each unit increase in BMI [75]. The results of our modeling appear in Table 8 below.
@@
Tables 6 and 7 of the preceding section establish the expectation that for a “typical male” 18–25 (having an “average comorbidities level”), prior-infection status determines whether vaccination is beneficial. Consistent with this expectation, Table 8 above indicates that for those having BMI in the intermediate range 25–34, prior-infection status does determine whether vaccination is beneficial. However, for those with high BMI (35–44) in a high-infection-risk scenario, Table 8 indicates that vaccination becomes beneficial even for those with prior-infection protection.
3.3. Sensitivity Analyses
Perhaps the most important of our sensitivity analyses assesses how the inclusion of 1st-dose benefits in our Scenarios B–E changes the corresponding benefit–risk ratios. This analysis, carried out in Section S2 of Supplement S2, yields the following benefit–risk ratios for our more likely scenarios: 0.68 for B, 0.73 for D, and 0.96 for E. Estimating inputs for 1st-dose benefit modeling was especially challenging given the extremely limited data on 1st-dose benefits for Omicron available before 22 January 2022. Our having lower confidence in input estimates for 1st-dose modeling contributed to our decision to present the modeling only as a sensitivity analysis.
Also, in Section S2 of Supplement S2, we provide a systematic sensitivity analysis related to Scenario D, our most likely scenario. For each input, we hold all others constant at their Scenario D values, determining the extent of possible variation in the chosen input before equipoise is reached. For example, if we fix our model input values, except for $I_{r}$, at their Scenario D levels, then benefit–risk equipoise is reached when $I_{r}\approx 0.678$, a value nearly 50% greater than the Scenario D value of $0.456$. We also point out that the risk-exceeds-benefit conclusion of Scenario D is not sensitive to the value of $E_{pi}$: even if $E_{pi}=0$, the benefit–risk ratio of this scenario is $\approx 0.83$. The insensitivity of the risk-exceeds-benefit outcome of Scenario D to significant changes in inputs from their Scenario D values increases confidence in the outcome.
3.4. Model Validation
@@
In Supplement S2, we test our model using Ontario COVID-hospitalization data from the evaluation period of 1 January 2022–31 May 2022, showing its projections of the hospitalizations-prevented benefit of vaccination are more accurate than those of the FDA [13]. We use Ontario data for many reasons; e.g., unlike the CDC, Public Health Ontario chose to publish incidental COVID-hospitalization rates (starting 8 January 2022) [36]; see [Supplement S2, Section S3] for other reasons.
We test our model in two ways: (i) showing that it fits Ontario COVID-hospitalization data with inputs at plausible levels for Ontario over the evaluation period [Supplement S2, Section S8], and (ii) using Ontario COVID-hospitalization data to predict corresponding data for the U.S. [Supplement S2, Section S7].
@@
Based on COVID-hospitalization counts from Public Health Ontario as well as corresponding incidental-hospitalization rates, we estimate that vaccinating one million 18–25-year-old Ontarians prevented 152 COVID hospitalizations over the evaluation period—see Table S2.5 of Supplement S2. Making various adjustments, e.g., accounting for the better health of Ontarians vs. Americans, we obtain an upper-bound estimate that 214 evaluation-period hospitalizations were prevented by one million full COVID vaccinations of 18–25–year-old males in the U.S.—see Table S2.7 of Supplement S2 and the discussion that follows it. Observe that 214 is an upper bound for our projections of Scenarios B–D (see Table 5) and a little below our projection 232 of Scenario E, but all of the FDA’s projections, ranging from 635 to 4766 for Omicron-based scenarios, substantially exceed 214.
@@
Ontario data also supports the definitiveness of the results of Table 6 above given the male has no comorbidities, suggesting that 151 in 1 million is an overestimate of his evaluation-period COVID hospitalization risk [Supplement S2, Section S2.3].
@@ 4. Discussion
By accounting for prior-infection protection and using evidence-based model inputs, we find in our Scenarios B–E of Table 5 that risks of mRNA-1273 vaccination outweighed benefits for the general population of 18–25-year-old males, assuming hospitalizations are the principal endpoint of analysis. In particular, for our scenarios predicting plausible infection-level increases over the evaluation period (Scenarios B, D, and E), our model suggests that vaccination of 18–25-year-old males generated between 16% and 62% more hospitalizations for vaccine-attributable myo/pericarditis (VAM/P) alone compared to COVID hospitalizations prevented, or, discounting 1st-dose VAM/P risk, between 8% and 52% more hospitalizations. Recalling that the FDA found, “[T]he benefits of vaccination with the Moderna COVID-19 vaccine clearly outweigh the risks, even among the male adolescent population” [13, Section 4], we have demonstrated that the conclusions of a benefit–risk assessment for vaccination may be dramatically impacted by accounting for the benefits conferred by prior infection. We have also demonstrated in Tables 68 how the benefit–risk picture for vaccination may vary based on an individual’s prior-infection and comorbidity status. Finally, we have demonstrated how a benefit–risk assessment for a two-dose vaccination series may acknowledge potential biases relating to the inclusion or exclusion of 1st dose benefits or risks.
@@
Consistent with our model’s finding that for 18–25-year-old males in the U.S., risks of mRNA-1273 vaccination outweighed benefits relative to hospitalizations, we note that France suspended mRNA-1273 use on 15 October 2021, making it available again on 8 November 2021 but at half-dose and only for those over 30 years [76, p. 4]. Also observe our model’s finding that vaccination of 18–25-year-old males generated between 8% and 52% more hospitalizations for vaccine-attributable myo/pericarditis (VAM/P) compared to COVID hospitalizations prevented is consistent with our model-validation work discussed in Section 3.4: recall that we estimate that at most 214 evaluation-period hospitalizations were prevented by one million mRNA-1273 vaccinations of 18–25-year-old males in the U.S. [Supplement 2, Table S2.7], and note that our corresponding VAM/P-hospitalization projection of 268 is a little over 25% higher than 214.
@@
We emphasize that the cardiac risk of COVID infection is incorporated into our modeling because our estimates of COVID hospitalizations include those for COVID-related cardiac issues. We address here assertions such as “Importantly, the risk of myocarditis after any dose of COVID-19 vaccine, in any population, is substantially lower than the risk of myocarditis after SARS-CoV-2 infection,” which appeared, without supporting references, in an April 2024 editorial of the European Heart Journal [77]. Contrary to the editorial’s assertion, Patone et al. find that, “In men younger than 40 years old, the number of excess myocarditis events [leading to hospitalization] per million people was higher after a 2nd dose of mRNA-1273 than after a positive SARS-CoV-2 test (97 [95% CI, 91–99] versus 16 [95% CI, 12–18])” [6]. Karlstad et al. also find significantly higher risk of hospitalized myocarditis after dose 2: 183.9 per million 2nd mRNA-1273 doses [5, Table 2] vs. 13.7 per million positive tests [5, eTable 7] among males 16–24. Moreover, these myocarditis-incidence rates expressed in cases per million positive tests overstate the risk of myocarditis after COVID infection expressed in cases per million COVID infections (because many infections are not confirmed by positive tests, but essentially all infections among  those seeking care for myocarditis are confirmed by positive tests) [78].  Consistent with [78], the authors of [79] note, “[T]he attributable risk estimates for COVID-19 [associated myocarditis] used laboratory confirmed cases as the denominator and will be affected by the proportion of all SARS-CoV-2 infections captured by testing, precluding a direct comparison with vaccine-associated attributable risks”.
4.1. The Strength of Prior-Infection Protection
@@
As noted in Section 1.3, at the time of the FDA’s mRNA-1273 assessment, immunological expectations and studies such as [21, 22, 23, 25, 32, 33] suggested that prior-infection protection would likely equal or exceed vaccine protection among the COVID-naïve. In our modeling, we have assumed that vaccination’s effectiveness against hospitalization is 77% in all scenarios, while that of prior infection is 88.5% except in Scenario E.
A systematic review [30] involving 21 vaccine-effectiveness studies from the period 3 December 2021–7 April 2022 reports,
@@
In contrast, the meta-analysis [27] estimates prior infection’s effectiveness against severe Omicron to be over 80% through 6 months after infection, while the meta-analysis [28, Table S2] estimates its effectiveness to be over 88% through 45 weeks after infection, with the former estimate based on studies from 1 January 2022–1 June 2022 and the latter, on studies published by 31 September 2022. In mid-January of 2022, the minimum reasonable evidence-based assumption would have been that prior-infection protection would equal vaccine protection (as in our Scenario E).
4.2. Assessing VAM/P Risk for mRNA-1273
Contributing to our finding mRNA-1273 vaccination’s risks exceeded benefits for 18–25-year-old males relative to hospitalizations is our assessment that VAM/P risk for mRNA-1273 is substantially greater than that suggested by the FDA’s BEST System. As detailed in Sections 1.4 and 2.3, our mRNA-1273 VAM/P-risk estimate of 268 hospitalizations per million full vaccinations is based on data, available to the FDA at the time of its mRNA-1273 assessment, from Ontario [14], England [15], and the United States [16, 12, 44, 51]. In Appendix S2 of Supplement S1, we pool data, unavailable at the time of the FDA’s mRNA-1273 assessment, from Scandinavia [5], France [7], and British Columbia [8] to obtain a VAM/P-hospitalization rate between 240 and 277 per million full, homologous mRNA-1273 vaccinations among males 18–25. A possible explanation for the FDA’s underestimating VAM/P risk for mRNA-1273 is that its BEST System has “inherent limitations, such as small sample sizes and imperfect sensitivity of ICD-10 codes to identify these rare outcomes” [12, Section 4].
Recall that there is evidence that the CDC’s VSD system also underascertained VAM/P risk for mRNA-1273. As noted in Section 1.5, during the period 1 June 2021–4 September 2021 of “enhanced passive surveillance” among males and females 18–39, Public Health Ontario’s myo/pericarditis detection rate per million 2nd doses of mRNA-1273 was 3.76 times that of the CDC’s VSD system based on its data through 9 October 2021 [Supplement S2, Appendix S3]. The study [4] offers at least a partial explanation for the VSD’s lower detection rate by revealing shortcomings in the VSD’s search algorithm.